OpenStudy (anonymous):

20 % of 2 is equal to @michele_laino @Nnesha

OpenStudy (alexandervonhumboldt2):

2*20/100=

OpenStudy (alexandervonhumboldt2):

multiply 2 by 20%. 20%=20/100=1/5. 2*1/5=?

OpenStudy (anonymous):

10?

OpenStudy (alexandervonhumboldt2):

no . 2*1/5=2/5

OpenStudy (anonymous):

oh okay

OpenStudy (anonymous):

thx @AlexandervonHumboldt2

OpenStudy (alexandervonhumboldt2):

np